# Xinyi Wang, Smaranda Tasmoc, Nantheera Anantrasirichai, and Angeliki Katsenou

### v1

@@ Abstract
Index Terms:
state space model, image banding, restoration, compression.
@@ I Introduction
Despite the outstanding performance of recent image and video coding standards, such as Versatile Video Coding (VVC) [1], Alliance for Open Media Video 1 (AV1) [2], and High Efficiency Video Coding (HEVC)/H.265 [3], streaming content live and/or at low bitrates requires aggressive compression, which often introduces compression artifacts that compromise visual quality. Banding is a type of quantization artifact that appears when a low-texture region of an image is encoded with insufficient bit depth. It is typically more noticeable in plain, smooth backgrounds, such as skylines, and becomes visible along a diagonal or curved edge when one color gradation shifts to another, creating a staircase-like effect. Banding is often encountered in user-generated content, as multiple rounds of transcoding on sharing platforms can further amplify the artifacts.
@@
Various techniques are designed for banding and contour artifacts in images and video based on the point of intervention, either before compression [4, 5, 6] or within an image/video encoder [7, 8] (known as in-loop processing). Here, we focus on post-processing methods because they are flexible and codec-independent. These methods either detect banded regions and intervene locally [9, 8, 10], or directly enhance the entire decoded image [11].
@@
Post-processing methods can be further categorized into traditional image processing techniques and deep learning based solutions. Examples of the first category include the following. Choi et al. [9] proposed a method combining directional dilation and edge-preserving filtering, where directional dilation enhances true edges while reducing false contours, and filtering protects image details from blurring. Yoo et al. [8] addressed contouring through an acyclic method that adds pseudo-random noise to blocks containing false contours, although it struggles with subtle artifacts. Bhagavathy et al. [11] introduced multiscale probabilistic dithering, which processes images at multiple scales using Gaussian pyramids and applies dithering at each level to mitigate contouring, though this can introduce noise and increase computational demands. Jin et al. [12] proposed a block-based detection approach using a composite model that balances gradient cues and blocks edge smoothness to eliminate false contours. Tu et al. [13] developed AdaDeband, an adaptive method with three key components: a banding detector, a content-aware filter that reconstructs gradients while preserving textures, and a dithering step that improves visual quality by re-quantizing smooth regions. Last, FFmpeg [14] provides two practical filters: Gradfun, which uses filtering and dithering on gradient areas to reduce banding, and deband, which focuses on removing pronounced banding while preserving detail. Both are less effective for heavily quantized inputs.
@@
Recent examples of deep learning-based solutions for banding include the deepDeband model [15], which employs a conditional Generative Adversarial Network (GAN) and is available in two variants: deepDeband-F, which is faster but less precise, and deepDeband-W, which adopts a bilateral weighting strategy to eliminate block discontinuity. In the same category of generative approaches falls cGAN [16] that uses a U-Net to first segment the image into a flat area mask, which then guides the generative network. Finally, the authors in [10] propose a method in the frequency domain that effectively isolates and eliminates components associated with banding artifacts, employing partitioned dynamic convolution for spatial differentiation of banded and non-banded features.
@@
In this paper, we present a debanding framework based on the alignment of a joint spatial-frequency representation with the human visual system. The frequency domain not only accurately captures the overall activity level of the luma component in an image, but also intuitively quantifies other perceptual features, such as contrast information. In banded areas, we observe medium to high frequencies in smooth areas, which opposes the pristine version. Based on this mechanism, we adopt a spatial frequency masking strategy. To this end, we employ the Wavelet State Space Model (WaveMamba) [17] and explore different strategies to create weighted wavelet maps, combining the benefits of both banded and restored images: preserving true edges, structures, and textures, while removing banding artifacts. In summary, this paper offers the following contributions:
@@ II Proposed Method
We employ the WaveMamba [17] architecture designed for low-light image enhancement and improve it by introducing a weighted wavelet map mechanism.
@@ II-A WaveMamba Network
WaveMamba [17] combines Discrete Wavelet Transform (DWT) and state-space modeling to capture low- and high-frequency areas, enabling simultaneous brightness and texture restoration. As banding artifacts mainly occur in low-frequency regions (smooth areas) and high-frequency details still need to be preserved in non-banded areas, WaveMamba can effectively separate low- and high-frequency information across multiple scales, suppressing artifacts and restoring textures.
@@
The overall WaveMamba network is constructed using a multi-scale U-Net [18] structure. After feature extraction and enhancement through the LFSSB and HFEB, the decoder reconstructs features in reverse. At each scale, high-frequency and low-frequency components are fused and refined using the same blocks, then upsampled via the Inverse Discrete Wavelet Transform (IDWT):
@@ III-A Datasets
There are not many public dataset available on debanding. The deepDeband dataset [15] consists of 1,439 pairs of ground truth and quantized full HD images (1920×1080) with annotated banded and non-banded regions, from which overlapping 256×256 patches were extracted. Only patches containing banding, along with their corresponding pristine counterparts, were retained, yielding 51,420 pairs of image patches. The dataset spans diverse visual content and is of sufficient scale to support the development of deep learning models. We trained our WaveMamba variants on deepDeband using a 70-20-10 split: 35,994 patch pairs for training, 10,284 for validation, and 5,142 for testing.
The BAND-2k [19] is the largest image dataset for evaluating banding artifacts, containing 2,000 distorted images generated using different compression and quantization methods. These images were sourced from over 870 diverse videos spanning computer graphics, user-generated, and professional content. BAND-2k includes subjective scores, along with patch-level annotations of banded and non-banded regions in quantized images. However, it is unsuitable for training models due to the lack of paired pristine and distorted image patches. Therefore, we use BAND-2k only for cross-dataset evaluation and inference in our experiments.
@@ III-B Evaluation Metrics
Objective quality metrics are essential for assessing the perceptual quality of enhanced images compared to their pristine counterparts. While widely adopted artifact-specific metrics such as Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [20], Learned Perceptual Image Patch Similarity (LPIPS) [21], and Video Multi-Method Assessment Fusion (VMAF) [22] have proven effective for general distortions, they fall short in detecting banding artifacts due to the subtle and localized nature of false contours.
To address this limitation, dedicated banding detection metrics such as Blind BANding Detector (BBAND) [23], Contrast-aware Multiscale Banding Index (CAMBI) [24], and Deep Banding Index (DBI) [25] have been proposed to quantify banding specifically. BBAND [23] is a video quality metric that estimates banding severity directly from the input. It computes a frame-level score based on the band visibility map and local gradient statistics, without requiring ground truth. CAMBI [24] improves perceptual alignment by incorporating the contrast sensitivity function [26] to capture human visual sensitivity. DBI [25] quantifies banding artifacts through a no-reference deep neural network model that produces both an overall banding score and a spatial banding map, accurately reflecting the severity and distribution of banding. Banding-specific metrics are expected to align better with human perception, although they may be less reliable under resolution changes or in non-banded regions.
@@ III-C Benchmarking Methods
To assess banding restoration, we benchmarked traditional and deep learning-based methods, ranging from classical signal processing filters to deep generative models. The Gradfun filter [14] in FFmpeg is a traditional technique for mitigating banding artifacts via gradient interpolation and dithering. It smooths pixel transitions in flat regions but fails to remove false contours in textured areas. The Plug-and-Play Priors (PPP) [27] framework incorporates learned denoising models as implicit priors, embedded within an optimization based on the alternating direction method of multipliers.
Deep Image Prior (DIP) [28] uses the implicit bias of convolutional networks as a handcrafted prior. An initialized network is gradually optimized to reconstruct a single corrupted image without external training data. For debanding, we use a randomly initialized U-Net to suppress banding artifacts during reconstruction. The Latent Diffusion Model (LDM) [29] is a generative model that operates in a perceptual latent space. It compresses an input image into a latent representation using a trained encoder, applies a diffusion process in the latent space, and reconstructs the image with a decoder. We adapt LDM for image super-resolution, treating banding removal as an enhancement of a degraded input. The deepDeband [15] is a GAN-based model built on the Pix2Pix framework, featuring a generator that produces debanded outputs and a discriminator that ensures realism. We include only deepDeband-F for full-image processing as the inference model in the performance comparison. To evaluate the effectiveness of the proposed frequency masking map, we integrate it into both deepDeband and our WaveMamba variants. We enhance deepDeband-F by fusing its output with the banded input via WWM, resulting in deepDeband-F-WWM, which helps suppress over-smoothing introduced by the generator. In WaveMamba variants, frequency masking is used to mitigate banding in smooth regions while enhancing fidelity in detailed areas during reconstruction.
@@ IV-A Performance Comparison
TABLE I: Quantitative comparison on the deepDeband [15] test set (patch pairs). Best and second-best results are highlighted in bold and underline, respectively. Arrows indicate if higher or lower values are preferred.
@@
FFmpeg [14] 31.383 0.941 0.110 0.955 0.632 0.035
PPP [27] 32.432 0.927 0.188 3.619 1.230 0.005
DIP [28] 34.364 0.957 0.093 1.728 0.820 0.034
LDM [29] 30.597 0.905 0.140 0.824 0.853 0.057
deepDeband-F [15] 34.584 0.964 0.071 0.057 0.464 0.122
@@
TABLE II: Cross-dataset evaluation of WaveMamba variants on BAND-2k [19] dataset, using models trained on deepDeband [15].
@@
Table I reports the quantitative comparison on the deepDeband test set. While metrics like CAMBI and BBAND are designed to assess banding artifacts, the results show that reduced banding does not lead to better structural consistency. For example, deepDeband-F and deepDeband-F-WWM achieve the lowest BBAND score (0.464) and the lowest CAMBI score (0.055), respectively, yet fail to improve PSNR or SSIM, indicating a lack of structural fidelity. In contrast, WaveMamba achieves the highest PSNR (42.399dB) but has the highest CAMBI and BBAND scores, indicating the worst banding artifacts, which do not align with visual inspection. Notably, the proposed DBI metric is well correlated with the Mean Opinion Score obtained from subjective quality assessment [25]. While PPP has the lowest DBI score (0.005), Fig. 2 shows that the debanded image appears blurry with noticeable banding transitions. WaveMamba and its variants achieve low DBI scores, consistent with visual inspection, indicating the closest alignment with human perception. The WaveMamba variants demonstrate that the proposed frequency masking map leads to lower DBI scores and achieves high PSNR and SSIM, proving its effectiveness in improving the removal of banding artifacts while maintaining structural consistency.

### v2

@@ I Introduction
Despite the outstanding performance of recent image and video coding standards, such as Versatile Video Coding (VVC) [VVC], Alliance for Open Media Video 1 (AV1) [AV1], and High Efficiency Video Coding (HEVC)/H.265 [HEVC], streaming content live and/or at low bitrates requires aggressive compression, which often introduces compression artifacts that compromise visual quality. Banding is a type of quantization artifact that appears when a low-texture region of an image is encoded with insufficient bit depth. It is typically more noticeable in plain, smooth backgrounds, such as skylines, and becomes visible along a diagonal or curved edge when one color gradation shifts to another, creating a staircase-like effect. Banding is often encountered in user-generated content, as multiple rounds of transcoding on sharing platforms can further amplify the artifacts.
@@
Various techniques are designed for banding and contour artifacts in images and video based on the point of intervention, either before compression [daly2003bit, 10.1145/383259.383326, 1057702] or within an image/video encoder [casali2015adaptive, yoo2009loop] (known as in-loop processing). Here, we focus on post-processing methods because they are flexible and codec-independent. These methods either detect banded regions and intervene locally [choi2006false, yoo2009loop, xu2024adaptive], or directly enhance the entire decoded image [bhagavathy2009multiscale].
@@
Post-processing methods can be further categorized into traditional image processing techniques and deep learning based solutions. Examples of the first category include the following. Choi et al. [choi2006false] proposed a method combining directional dilation and edge-preserving filtering, where directional dilation enhances true edges while reducing false contours, and filtering protects image details from blurring. Yoo et al. [yoo2009loop] addressed contouring through an acyclic method that adds pseudo-random noise to blocks containing false contours, although it struggles with subtle artifacts. Bhagavathy et al. [bhagavathy2009multiscale] introduced multiscale probabilistic dithering, which processes images at multiple scales using Gaussian pyramids and applies dithering at each level to mitigate contouring, though this can introduce noise and increase computational demands. Jin et al. [jin2011composite] proposed a block-based detection approach using a composite model that balances gradient cues and blocks edge smoothness to eliminate false contours. Tu et al. [tu2020adaptive] developed AdaDeband, an adaptive method with three key components: a banding detector, a content-aware filter that reconstructs gradients while preserving textures, and a dithering step that improves visual quality by re-quantizing smooth regions. Last, FFmpeg [ffmpeg] provides two practical filters: Gradfun, which uses filtering and dithering on gradient areas to reduce banding, and deband, which focuses on removing pronounced banding while preserving detail. Both are less effective for heavily quantized inputs.
@@
Recent examples of deep learning-based solutions for banding include the deepDeband model [zhou2022deep], which employs a conditional Generative Adversarial Network (GAN) and is available in two variants: deepDeband-F, which is faster but less precise, and deepDeband-W, which adopts a bilateral weighting strategy to eliminate block discontinuity. In the same category of generative approaches falls cGAN [chang2025two] that uses a U-Net to first segment the image into a flat area mask, which then guides the generative network. Finally, the authors in [xu2024adaptive] propose a method in the frequency domain that effectively isolates and eliminates components associated with banding artifacts, employing partitioned dynamic convolution for spatial differentiation of banded and non-banded features.
@@
In this paper, we present a debanding framework based on the alignment of a joint spatial-frequency representation with the human visual system. The frequency domain not only accurately captures the overall activity level of the luma component in an image, but also intuitively quantifies other perceptual features, such as contrast information. In banded areas, we observe medium to high frequencies in smooth areas, which opposes the pristine version. Based on this mechanism, we adopt a spatial frequency masking strategy. To this end, we employ the Wavelet State Space Model (WaveMamba) [zou2024wave] and explore different strategies to create weighted wavelet maps, combining the benefits of both banded and restored images: preserving true edges, structures, and textures, while removing banding artifacts. In summary, this paper offers the following contributions:
@@ II Proposed Method
We employ the WaveMamba [zou2024wave] architecture designed for low-light image enhancement and improve it by introducing a weighted wavelet map mechanism.
@@ II-A WaveMamba Network
WaveMamba [zou2024wave] combines Discrete Wavelet Transform (DWT) and state-space modeling to capture low- and high-frequency areas, enabling simultaneous brightness and texture restoration. As banding artifacts mainly occur in low-frequency regions (smooth areas) and high-frequency details still need to be preserved in non-banded areas, WaveMamba can effectively separate low- and high-frequency information across multiple scales, suppressing artifacts and restoring textures.
@@
The overall WaveMamba network is constructed using a multi-scale U-Net [ronneberger2015u] structure. After feature extraction and enhancement through the LFSSB and HFEB, the decoder reconstructs features in reverse. At each scale, high-frequency and low-frequency components are fused and refined using the same blocks, then upsampled via the Inverse Discrete Wavelet Transform (IDWT):
@@ III-A Datasets
There are not many public dataset available on debanding. The deepDeband dataset [zhou2022deep] consists of 1,439 pairs of ground truth and quantized full HD images (1920×1080) with annotated banded and non-banded regions, from which overlapping 256×256 patches were extracted. Only patches containing banding, along with their corresponding pristine counterparts, were retained, yielding 51,420 pairs of image patches. The dataset spans diverse visual content and is of sufficient scale to support the development of deep learning models. We trained our WaveMamba variants on deepDeband using a 70-20-10 split: 35,994 patch pairs for training, 10,284 for validation, and 5,142 for testing.
The BAND-2k [chen2024band] is the largest image dataset for evaluating banding artifacts, containing 2,000 distorted images generated using different compression and quantization methods. These images were sourced from over 870 diverse videos spanning computer graphics, user-generated, and professional content. BAND-2k includes subjective scores, along with patch-level annotations of banded and non-banded regions in quantized images. However, it is unsuitable for training models due to the lack of paired pristine and distorted image patches. Therefore, we use BAND-2k only for cross-dataset evaluation and inference in our experiments.
@@ III-B Evaluation Metrics
Objective quality metrics are essential for assessing the perceptual quality of enhanced images compared to their pristine counterparts. While widely adopted artifact-specific metrics such as Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [wang2004image], Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable], and Video Multi-Method Assessment Fusion (VMAF) [li2016toward] have proven effective for general distortions, they fall short in detecting banding artifacts due to the subtle and localized nature of false contours.
To address this limitation, dedicated banding detection metrics such as Blind BANding Detector (BBAND) [tu2020bband], Contrast-aware Multiscale Banding Index (CAMBI) [tandon2021cambi], and Deep Banding Index (DBI) [kapoor2021capturing] have been proposed to quantify banding specifically. BBAND [tu2020bband] is a video quality metric that estimates banding severity directly from the input. It computes a frame-level score based on the band visibility map and local gradient statistics, without requiring ground truth. CAMBI [tandon2021cambi] improves perceptual alignment by incorporating the contrast sensitivity function [seshadrinathan2009image] to capture human visual sensitivity. DBI [kapoor2021capturing] quantifies banding artifacts through a no-reference deep neural network model that produces both an overall banding score and a spatial banding map, accurately reflecting the severity and distribution of banding. Banding-specific metrics are expected to align better with human perception, although they may be less reliable under resolution changes or in non-banded regions.
@@ III-C Benchmarking Methods
To assess banding restoration, we benchmarked traditional and deep learning-based methods, ranging from classical signal processing filters to deep generative models. The Gradfun filter [ffmpeg] in FFmpeg is a traditional technique for mitigating banding artifacts via gradient interpolation and dithering. It smooths pixel transitions in flat regions but fails to remove false contours in textured areas. The Plug-and-Play Priors (PPP) [venkatakrishnan2013plug] framework incorporates learned denoising models as implicit priors, embedded within an optimization based on the alternating direction method of multipliers.
Deep Image Prior (DIP) [ulyanov2018deep] uses the implicit bias of convolutional networks as a handcrafted prior. An initialized network is gradually optimized to reconstruct a single corrupted image without external training data. For debanding, we use a randomly initialized U-Net to suppress banding artifacts during reconstruction. The Latent Diffusion Model (LDM) [rombach2022high] is a generative model that operates in a perceptual latent space. It compresses an input image into a latent representation using a trained encoder, applies a diffusion process in the latent space, and reconstructs the image with a decoder. We adapt LDM for image super-resolution, treating banding removal as an enhancement of a degraded input. The deepDeband [zhou2022deep] is a GAN-based model built on the Pix2Pix framework, featuring a generator that produces debanded outputs and a discriminator that ensures realism. We include only deepDeband-F for full-image processing as the inference model in the performance comparison. To evaluate the effectiveness of the proposed frequency masking map, we integrate it into both deepDeband and our WaveMamba variants. We enhance deepDeband-F by fusing its output with the banded input via WWM, resulting in deepDeband-F-WWM, which helps suppress over-smoothing introduced by the generator. In WaveMamba variants, frequency masking is used to mitigate banding in smooth regions while enhancing fidelity in detailed areas during reconstruction.
@@ IV-A Performance Comparison
TABLE I: Quantitative comparison on the deepDeband [zhou2022deep] test set (patch pairs). Best and second-best results are highlighted in bold and underline, respectively. Arrows indicate if higher or lower values are preferred.
@@
FFmpeg [ffmpeg] 31.383 0.941 0.110 0.955 0.632 0.035
PPP [venkatakrishnan2013plug] 32.432 0.927 0.188 3.619 1.230 0.005
DIP [ulyanov2018deep] 34.364 0.957 0.093 1.728 0.820 0.034
LDM [rombach2022high] 30.597 0.905 0.140 0.824 0.853 0.057
deepDeband-F [zhou2022deep] 34.584 0.964 0.071 0.057 0.464 0.122
@@
TABLE II: Cross-dataset evaluation of WaveMamba variants on BAND-2k [chen2024band] dataset, using models trained on deepDeband [zhou2022deep].
@@
Table I reports the quantitative comparison on the deepDeband test set. While metrics like CAMBI and BBAND are designed to assess banding artifacts, the results show that reduced banding does not lead to better structural consistency. For example, deepDeband-F and deepDeband-F-WWM achieve the lowest BBAND score (0.464) and the lowest CAMBI score (0.055), respectively, yet fail to improve PSNR or SSIM, indicating a lack of structural fidelity. In contrast, WaveMamba achieves the highest PSNR (42.399dB) but has the highest CAMBI and BBAND scores, indicating the worst banding artifacts, which do not align with visual inspection. Notably, the proposed DBI metric is well correlated with the Mean Opinion Score obtained from subjective quality assessment [kapoor2021capturing]. While PPP has the lowest DBI score (0.005), Fig. 2 shows that the debanded image appears blurry with noticeable banding transitions. WaveMamba and its variants achieve low DBI scores, consistent with visual inspection, indicating the closest alignment with human perception. The WaveMamba variants demonstrate that the proposed frequency masking map leads to lower DBI scores and achieves high PSNR and SSIM, proving its effectiveness in improving the removal of banding artifacts while maintaining structural consistency.